# Ultra-massive spacetimes in 2+1 dimensions with positive $\Lambda$

### v1

@@ 2 Some expressions for 3-dimensional (locally) de Sitter spacetime
$\partial_{z}=X\partial_{Y}-Y\partial_{X}\ ,\hskip 22.76219pt\partial_{\varphi}=-U\partial_{Z}-Z\partial_{U}\ .$ (12)
@@ 3 The models
and then $m(v)$ is called the mass function. The metric (13) is a solution of the Einstein field equations
@@
Notice that as a one-form this is just $dr$. Thus, computing the norm of $\vec{H}$ we deduce that the round circles are future-trapped if $r^{2}>F(v)/\Lambda$, and marginally future-trapped if $r^{2}=F(v)/\Lambda$. Hence, they can only exist in regions with $F(v)>0$. In these regions, there is a surface foliated by these marginally trapped circles, that is to say, a circularly symmetric MTT, defined by
@@ 4 Discussion
Such an absence of dynamical horizons is due to the instability in spatial directions of all possible marginally trapped circles. In what follows we provide a transparent argument showing that there can certainly be MTTs —as shown in Figs.25—, but they will never form a dynamical horizon. Let us consider the stability operator for marginally trapped co-dimension 2 submanifolds, as introduced in [2, 3], see also section 2 in [23]. The main reason why the ultra-massive cases arise in 4 dimensions is the upper bound for the area of spatially stable marginally trapped spheres, and ultimately this bound appears due to the stability property. This is clear from the application of the stability operator, which leads to formula (16) in [23]
@@
where one has used the Gauss-Bonnet theorem when integrating the scalar curvature on the marginally trapped sphere $S$. Here, $A_{S}$ is the area of the sphere, $k^{\mu}$ is the null vector normal to $S$ with vanishing expansion–i.e, pointing along the mean curvature vector–, $\ell^{\mu}$ the other null vector field normal to $S$ (normalized $\ell^{\mu}k_{\mu}=-1$), $n^{\mu}$ is the vector field normal to $S$ along which the variation of the null expansion is performed (normalized $k^{\mu}n_{\mu}=1$), $W$ is a non-negative quantity formed from $T_{\mu\nu}$ and the shear tensor of the surface, and $\lambda_{n}$ is the principal eigenvalue of the stability operator ruling the variation of the vanishing expansion in the $\vec{n}$-direction [2, 3]. A non-negative $\lambda_{n}$ means stability along the direction $n^{\mu}$. It must be noticed that, if DEC holds and $n^{\mu}$ is spacelike, all the terms in the integrand are non-negative and thus $(\Lambda+\lambda_{n})\leq 4\pi$ leading to the area bound for marginally trapped spheres that are stable in the spatial direction $n^{\mu}$.
@@
where $L_{\cal C}$ is the length of ${\cal C}$. In this case, if $\Lambda>0$ and DEC holds, then $\lambda_{n}<0$ is necessarily negative for any spacelike $n^{\mu}$. In other words, all marginally trapped circles are unstable in spatial directions. Therefore, marginally trapped circles perturbed outwardly within a spacelike slice inevitably lead to trapped circles. This implies that any MTT cannot be spacelike and, a fortiori, leads to the impossibility of having an event horizon for future null infinity. Thus, only conical (and naked) singularities may arise if future null infinity remains. If not, the ultra-massive cases (and the critical case) do emerge, and turn out to provide a natural ‘home’ for the locally de Sitter spacetimes found by Mess.

### v2

@@ 2 Some expressions for 3-dimensional (locally) de Sitter spacetime
$\partial_{z}=X\partial_{Y}-Y\partial_{X}\ ,\hskip 22.76219pt\partial_{\varphi}=U\partial_{Z}+Z\partial_{U}\ .$ (12)
@@ 3 The models
and then $m(v)$ is called the mass function. It equals zero in de Sitter space, and will be non-negative in all the solutions we consider. The metric (13) is a solution of the Einstein field equations
@@
Notice that as a one-form this is just $dr$. Thus, computing the norm of $\vec{H}$ we deduce that the round circles are future-trapped if $r^{2}<F(v)/\Lambda$, and marginally future-trapped if $r^{2}=F(v)/\Lambda$. Hence, they can only exist in regions with $F(v)>0$. In these regions, there is a surface foliated by these marginally trapped circles, that is to say, a circularly symmetric MTT, defined by
@@ 4 Discussion
Such an absence of dynamical horizons is due to the instability in spatial directions of all possible marginally trapped circles. In what follows we provide a transparent argument showing that there can certainly be MTTs —as shown in Figs.25—, but they will never form a dynamical horizon. Let us consider the stability operator for marginally trapped co-dimension 2 submanifolds, as introduced in [2, 3], see also section 2 in [23]. The main reason why the ultra-massive cases arise in 4 dimensions is the upper bound for the area of spatially stable marginally trapped spheres, and ultimately this bound appears due to the stability property. This is clear from the application of the stability operator, as we explain next.
Any co-dimension two marginally trapped submanifold $S$ has a vanishing expansion $\theta_{k}$ along one normal null direction $k^{\mu}$, meaning that $\theta_{k}:=H_{\mu}k^{\mu}=0$. However, if we perturb the submanifold along an arbitrary vector field $n^{\mu}$ normal to $S$ (normalized $k^{\mu}n_{\mu}=1$), the corresponding perturbed mean curvature vector field loses (in general) its null character and acquires, at first order, a non-zero expansion given by
$\delta_{f\vec{n}}\theta^{k}:=L_{n}f$ (20)
where $f$ is a function on $S$ providing the “amount” of the perturbation along $\vec{n}$ at each point on $S$. $L_{n}$ is called the stability operator for $S$ in the direction $\vec{n}$ and is explicitly given by [2, 3]
$L_{n}f=-\Delta f+2s^{B}D_{B}f+f\left(\frac{\bar{R}}{2}-s^{B}s_{B}+D_{B}s^{B}-G_{\mu\nu}k^{\mu}\ell^{\nu}-\frac{n^{\rho}n_{\rho}}{2}\,W\right)$ (21)
where $\bar{R}$ is the scalar curvature on $S$, $D_{B}$ its covariant derivative, $\Delta=D_{B}D^{B}$ its Laplacian, $\ell^{\mu}$ the other independent null vector field normal to $S$ (normalized $\ell^{\mu}k_{\mu}=-1$), $s_{B}$ the one-form on $S$ defined by
[…]
and finally
$W:=G_{\mu\nu}k^{\mu}k^{\nu}+\Sigma_{k}^{2}$
with $\Sigma_{k}$ the shear scalar of $k^{\mu}$. Observe that, if the Einstein field equations and the null convergence condition hold, $W$ is non-negative.
$L_{n}$ is a (generally non-self-adjoint) elliptic operator on $S$. Nevertheless, as proven in [2, 3], it possesses a real principal eigenvalue denoted by $\lambda_{n}$. The corresponding real eigenfunction, denoted by $\phi_{n}$, can be chosen to be positive on all of $S$. A non-negative $\lambda_{n}$ means stability of the marginally trapped $S$ along the direction $n^{\mu}$.
For positive $f>0$, an equivalent way of writing (21) is [24, 23]
$\frac{L_{n}f}{f}=-\Delta\ln f+D_{B}s^{B}-\left(D_{B}\ln f-s_{B}\right)\left(D^{B}\ln f-s^{B}\right)+\frac{\bar{R}}{2}-\frac{n^{\rho}n_{\rho}}{2}W-G_{\mu\nu}k^{\mu}\ell^{\nu}$ (22)
which, applied to $f=\phi_{n}$ and using $L_{n}\phi_{n}=\lambda_{n}\phi_{n}$, leads in 3+1 dimensions to formula (16) in [23]
@@
where one has used the Gauss-Bonnet theorem when integrating the scalar curvature on the marginally trapped sphere $S$. Here, $A_{S}$ denotes the area of the sphere. It must be noticed that, if DEC holds and $n^{\mu}$ is spacelike, all the terms in the integrand are non-negative and thus $(\Lambda+\lambda_{n})\leq 4\pi$ leading to the area bound for marginally trapped spheres that are stable in the spatial direction $n^{\mu}$.
@@
where now $L_{\cal C}$ is the length of ${\cal C}$. In this case, if $\Lambda>0$ and DEC holds, then $\lambda_{n}<0$ is necessarily negative for any spacelike $n^{\mu}$. In other words, all marginally trapped circles are unstable in spatial directions. Therefore, marginally trapped circles perturbed outwardly within a spacelike slice inevitably lead to trapped circles. This implies that any MTT cannot be spacelike and, a fortiori, leads to the impossibility of having an event horizon for future null infinity. Thus, only conical (and naked) singularities may arise if future null infinity remains. If not, the ultra-massive cases (and the critical case) do emerge, and turn out to provide a natural ‘home’ for the locally de Sitter spacetimes found by Mess.